A 1.92 g sample of a soluble salt dissolves in 22.9 mL of water in a coffee cup calorimeter. When the salt dissolves the temperature in the calorimeter decreases by 11.6 °C. What is the energy per gram (J/g) for the dissolution of the salt?

Assume the specific heat of the solution is the same as water; 4.184 J/g°C, that the density of water is 1.00 g/mL, and that the calorimeter itself does not absorb or release any heat.
1
Expert's answer
2020-10-04T09:15:33-0400

Water

"Mass=22.9ml\newline
Specific\ heat\ capacity=4.184J/g^0C\newline \Delta T=11.6°C"

Soluble salt

"Mass=1.92 g"

Heat lose of water is the same as heat lose of solution since there

is no heat exchange.

"(\Delta H)_{sol}\ or (q)= ms\Delta T"

"m= mass\ of\ solution \newline m=22.9+1.92=24.82g \newline s=4.184J/g°C\newline \Delta T=-11.6°C\ since\ it's\ a\ decrease"

Plugging in the values into the formula;

"q= 24.82*4.184*-11.6\newline q=-1204.6J"

For one mole of the salt

"-1204.6\u00f71= -1204.6J\/g"